The National Sleep Foundation used a survey to determine whether hours of sleeping per night are independent of age (Newsweek, January 19, 2004). The following show the hours of sleep on weeknights for a sample of individuals age 49 and younger and for a sample of individuals age 50 and older Age 49 or younger 50 or older a. Conduct a test of independence to determine whether the hours of sleep on weeknights are independent of age. Use α = .05. Use Table 12.4 Fewer than 6 36 38 Hours of Sleep 6 to 6.9 57 56 7 to 7.9 76 72 8 or more Total 240 260 71 94 Compute the value of the χ 2 test statistic (to 2 decimals) The p value is greater than .10 What is your conclusion? Cannot reject the assumption that age and hours of sleep are independent ▼ b. Using the total sample of 500, estimate the percentage of people who sleep less than 6, 6 to 6.9, 7 to 7.9, and 8 or more hours on weeknights (to 1 decimal) Less than 6 hours 6 to 6.9 hours 7 to 7.9 hours 8 or more hours
Answer #1

(a)

(i)

Observed Frequencies:

Fewer than 6 6 to 6.9 7 to 7.9 8 or more Total
49 or younger 36 57 76 71 240
50 or older 38 56 72 94 260
Total 74 113 148 165 500

Expected Frequencies:

Fewer than 6 6 to 6.9 7 to 7.9 8 or more Total
49 or younger 74X240/500=35.52 54.24 71.04 79.20 240
50 or older 38.48 58.76 76.96 85.80 260
Total 74 113 148 165 500

The value of \chi ^{2} statistic is obtained as follows:

O E (O -E)2/E
36 35.52 0.01
57 54.24 0.14
76 71.04 0.35
71 79.20 0.85
38 38.48 0.01
56 58.76 0.13
72 76.96 0.32
94 85.80 0.78
Total = \chi ^{2} = 2.58

The value of the \chi ^{2} statistic is 2.58

(ii)

ndf = (r - 1) X (c - 1) = (2 - 1) X (3 - 1) = 2

By Technology, p - value = 0.4608

So,

Correct option:

The p - value is greater than .10

(iii)

Since p - value is greater than \alpha , the difference is not significant. Fail to reject null hypothesis.

So,

Correct option:

cannot reject the assumption that age and hours of sleep are independent.

(b)

Less than 6 hours = (74/500) X 100 = 14.8 %

6 to 6.9 hours =(113/500) X 100 = 22.6 %

7 to 7.9 hours = (148/500) X 500 = 29.6 %

8 or more hours = (165/500) X 100 = 33.0 %